OpenStudy (anonymous):

1000 is borrowed at a rate of 16% interest per annul. Find the amount due at the end of two years if the interest is compounded annually?

OpenStudy (anonymous):

I tried it, but my answer is wrong. I don't know what Im doing wrong tho.

OpenStudy (anonymous):

1000 x 0.16 = 160 1160 x 0.16 = 185.6 so 1160 + 185.6 answer = 1345.6

OpenStudy (tkhunny):

What did you do? What were the results of your trying?

OpenStudy (anonymous):

I have the formula A(t)= P(1+i/n)^nt were P is the initial amount, I is the interest rate, t is time, and n is the compound years.

OpenStudy (anonymous):

for simple interest like this just do what i do :) for more difficult use the formula

OpenStudy (anonymous):

A(2)=1000(1+1.16)^(0.16)(2)

OpenStudy (anonymous):

I have to use the formula because my teacher wants to see it, I have a test tomorrow on this and I don't think I can learn a new way now.

OpenStudy (tkhunny):

A(t)= P(1+i/n)^nt That's fine, n = 1 t =2 P = 1000

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first year is 1000 x 0.16 ( interest) giving 160 add 160 to 1000 to give end of year 1 amount 1160x 0.16 ( second year interest) = 185.6 add total year 1 with interest of year 2 1160+185.6 is = 1345.6

OpenStudy (tkhunny):

Yes, you can do it a year at a time if you like. This is very impractical for longer periods and shorter compounding periods. Let the algebra help you. \(1000(1 + 0.16)^{2} = 1345.6\)

OpenStudy (anonymous):

why is it to the power of two?

OpenStudy (anonymous):

instead of 0.32

OpenStudy (tkhunny):

I = 0.16 n = 1 t = 2 There is no "I" in the exponent.

OpenStudy (anonymous):

I know what I was doing wrong, I always mix up the i and n. Thank you

OpenStudy (tkhunny):

That will, indeed, produce an incorrect calculation. Good work!